OpenStudy (anonymous):

Graph the function by hand. Find the vertex; round approx coodinates to the second decimal place. 3x^2-y+5=6x

OpenStudy (anonymous):

Let's get it into y=_____ forms 3x^2-6x+5=y

OpenStudy (anonymous):

vertex equation -b ---- 2a

OpenStudy (anonymous):

Wasn't correct according to the computer, but thanks anyway.

OpenStudy (anonymous):

I am not there yet

OpenStudy (anonymous):

6 ---- = -1/2 -12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plug in -1/2 into function

OpenStudy (anonymous):

(-1/2,35/4)

OpenStudy (anonymous):

ax^2+bx+c=y : general equation 3x^2-6x+5=y : ur equation for vertex x=-b/(2a) for ur equation a=3 b=-6 c=5 is what imranmeah did